# Structurally damped semilinear evolution equation for positive operators on Hilbert space

### v1

@@ 6. Application
$\|\mathcal{N}u\|_{X^{k,\beta}}$$\leq$ $L,\quad\text{for some suitable }L>0,$ (6.2)
$\|\mathcal{N}u-\mathcal{N}v\|_{X^{k,\beta}}$$\leq$ $\frac{1}{r}\|u-v\|_{X^{k,\beta}},\quad\text{for some suitable }r>1,$ (6.3)
for all $u,v\in X^{k,\beta}$. We will establish that $\mathcal{N}$ is a contraction mapping on the Banach space $X^{k,\beta}$ using the inequalities (6.2) and (6.3). As a result, the Banach fixed-point theorem asserts that $\mathcal{N}$ admits a unique fixed point in $X^{k,\beta}$, which corresponds to a unique global solution $u$ of the equation $\mathcal{N}u=u$. We now proceed to prove the inequalities (6.2) and (6.3).
Estimate for $\|\mathcal{N}u\|_{X^{k,\beta}}$: Employing the definition of the norm $\|\cdot\|_{X^{k,\beta}}$ and the estimates provide in Theorem 2.8, we get the following:
$\|u^{\operatorname{lin}}(t)\|_{X^{k,\beta}}<\infty,$ (6.4)
and hence $u^{\operatorname{lin}}\in X^{k,\beta}$. To determine the norm $\left\|u^{\operatorname{non}}\right\|_{X^{k,\beta}}$, the following component norms must be estimated:
$\|\partial_{t}^{i}\mathcal{L}^{j}u^{\operatorname{non}}(t)\|_{\mathcal{H}},\quad(i,j)\in\{(0,0),(0,\beta),(k,0)\}.$
Using the estimates from Theorem 2.8 for $u_{0}=0$ and $u_{1}=f(u)$, we get
$\|\partial_{t}^{i}\mathcal{L}^{j}u^{\operatorname{non}}(t)\|_{\mathcal{H}}=\left\|\int_{0}^{t}E_{1}(t-\tau)*_{\mathcal{L}}f(u)(t)\mathrm{d}\tau\right\|_{\mathcal{H}}\leq\int_{0}^{t}e^{-\delta(t-\tau)}\left\|f(u)(\tau)\right\|_{\mathcal{H}_{\mathcal{L}}^{n_{i,j}}}\mathrm{d}\tau\\ \leq e^{-\delta t}\int_{0}^{t}e^{-\delta(p-1)\tau}e^{\delta p\tau}\left\|u(\tau)\right\|^{p}_{\mathcal{H}_{\mathcal{L}}^{n_{i,j}}}\mathrm{d}\tau\leq e^{-\delta t}\int_{0}^{t}e^{-\delta(p-1)\tau}\mathrm{d}\tau\|u\|^{p}_{X^{k,\beta}}\leq e^{-\delta t}\|u\|^{p}_{X^{k,\beta}}.$ (6.5)
@@
$\|\mathcal{N}u\|_{X^{k,\beta}}\leq A\|u_{0}\|_{\mathcal{H}_{\mathcal{L}}^{m_{k,\beta}}}+A\|u_{1}\|_{\mathcal{H}_{\mathcal{L}}^{n_{k,\beta}}}+A\|u\|_{X^{k,\beta}},$ (6.6)
@@
Estimate for $\|\mathcal{N}u-\mathcal{N}v\|_{X^{k,\beta}}$: For $u,v\in X^{k,\beta}$, the Definition 6.1 yields
@@
Verifying that $\|u^{\mathrm{lin}}-v^{\mathrm{lin}}\|_{X^{k,\beta}}=0$ is straightforward. We must now estimate the following norms in order to calculate the norm $\left\|u^{\operatorname{non}}-v^{\operatorname{non}}\right\|_{X^{k,\beta}}$:
$\|\partial_{t}^{i}\mathcal{L}^{j}(u^{\operatorname{non}}-v^{\operatorname{non}})(t)\|_{\mathcal{H}},\quad(i,j)\in\{(0,0),(0,\beta),(k,0)\}.$
Utilizing the estimate analogous to the estimate $\left\|\partial_{t}^{i}\mathcal{L}^{j}u^{\operatorname{non}}(t,\cdot)\right\|_{\mathcal{H}}$, we get
[…]
@@
$\left\|\partial_{t}^{i}\mathcal{L}^{j}\left(u^{\operatorname{non}}-v^{\operatorname{non}}\right)(t)\right\|_{\mathcal{H}}\lesssim\|u-v\|_{X^{k,\beta}}\left(\|u\|_{X^{k,\beta}}^{p-1}+\|v\|^{p-1}_{X^{k,\beta}}\right).$
Thus, combining the above estimate with $\|u^{\mathrm{lin}}-v^{\mathrm{lin}}\|_{X^{k,\beta}}=0$, we have
$\|\mathcal{N}u-\mathcal{N}v\|_{X^{k,\beta}}\leq B\|u-v\|_{X^{k,\beta}}\left(\|u\|_{X^{k,\beta}}^{p-1}+\|v\|^{p-1}_{X^{k,\beta}}\right),$ (6.7)
@@
Let us set $R:=rA(\|u_{0}\|_{\mathcal{H}_{\mathcal{L}}^{m_{k,\beta}}}+\|u_{1}\|_{\mathcal{H}_{\mathcal{L}}^{n_{k,\beta}}})$ for some $r>1$ with sufficiently small initial Cauchy data $\|(u_{0},u_{1})\|_{\mathcal{H}_{\mathcal{L}}^{m_{k,\beta}}\times\mathcal{H}_{\mathcal{L}}^{n_{k,\beta}}}<\varepsilon$ such that
@@
$\|\mathcal{N}u\|_{X^{k,\beta}}\leq\frac{2R}{r}\quad\text{and }\quad\|\mathcal{N}u-\mathcal{N}v\|_{X^{k,\beta}}\leq\frac{1}{r}\|u-v\|_{X^{k,\beta}},$
respectively for all $u,v\in\mathscr{B}_{R}:=\{u\in X^{k,\beta}:\|u\|_{X}\leq R\}$. As a closed ball of Banach space $X^{k,\beta}$, $\mathscr{B}_{R}$ is a Banach space itself. This allows us to conclude the proof of the theorem using the Banach fixed theorem. ∎
@@ 7. Final remarks
    The global (in-time) result indirectly gives the Fujita-type critical exponent based on the assumption (6.1), which is intrinsically reliant on the choice of the Hilbert space and the associated operator $\mathcal{L}$. The conclusion indicates a limitation, as the essential threshold is not explicitly stated but rather embedded in structural assumptions related to the functional context.
@@
    Since the decay estimates derived in the damped instances are generally weaker, they are frequently not enough to provide the time integrability needed to handle the nonlinear factors. Our investigation of proving the local-in-time existence of solutions in such cases will be addressed in the follow-up work of this article.

### v2

@@ 6. Application
$\|\mathcal{N}u\|_{X}$ $\leq$ $L,\quad\text{for some suitable }L>0,$ (6.7)
[…]
for all $u,v\in X$. We will establish that $\mathcal{N}$ is a contraction mapping on the Banach space $X$ using the inequalities (6.7) and (6.8). As a result, the Banach fixed-point theorem asserts that $\mathcal{N}$ admits a unique fixed point in $X$, which corresponds to a unique global solution $u$ of the equation $\mathcal{N}u=u$. We now proceed to prove the inequalities (6.7) and (6.8).
Estimate for $\|\mathcal{N}u\|_{X}$: Employing the definition of the norm $\|\cdot\|_{X}$ and the estimates provide in Theorem 2.5, we get the following:
$\|u^{\operatorname{lin}}(t)\|_{X}\lesssim\|(u_{0},u_{1})\|_{\mathcal{A}^{\sigma,\theta}},$ (6.9)
and hence $u^{\operatorname{lin}}\in X$. To determine the norm $\left\|u^{\operatorname{non}}\right\|_{X}$, the following component norms must be estimated:
$\|\partial_{t}^{i}\mathcal{L}^{j}u^{\operatorname{non}}(t)\|_{2},\quad(i,j)\in\{(0,0),(0,\sigma/2),(1,0)\}.$
Using the assumption (6.5) on $f$, we get
$\|(f(u)-f(v))(t,\cdot)\|_{2}^{2}\lesssim\int_{\Omega}\left(|u(t,\cdot)|^{p-1}+|v(t,\cdot)|^{p-1}\right)^{2}|u(t,\cdot)-v(t,\cdot)|^{2}.$
Consequently, by the Hölder inequality, we get
$\|(f(u)-f(v))(t,\cdot)\|_{2}^{2}\lesssim\left(\|u(t,\cdot)\|_{2p}^{p-1}+\|v(t,\cdot)\|_{2p}^{p-1}\right)^{2}\|(u-v)(t,\cdot)\|_{2p}^{2},$
since $\frac{1}{\frac{p}{p-1}}+\frac{1}{p}=1$. Now, by using the Gagliardo-Nirenberg inequality (6.1) and the Young’s inequality
$x^{\alpha}y^{1-\alpha}\leq\alpha x+(1-\alpha)y,\quad\text{for }0\leq\alpha\leq 1\quad\text{and }x,y\geq 0,$
we obtain
$\|(f(u)-f(v))(t,\cdot)\|_{2}$ $\lesssim\left[\left(\|\mathcal{L}^{\sigma/2}u(t,\cdot)\|^{\alpha}_{2}\|u\|_{2}^{1-\alpha}\right)^{p-1}+\left(\|\mathcal{L}^{\sigma/2}v(t,\cdot)\|^{\alpha}_{2}\|v(t,\cdot)\|_{2}^{1-\alpha}\right)^{p-1}\right]$
$\times\left(\|\mathcal{L}^{\sigma/2}(u-v)(t,\cdot)\|^{\alpha}_{2}\|(u-v)(t,\cdot)\|_{2}^{1-\alpha}\right)$
$\lesssim\left[\left(\left\|\mathcal{L}^{\sigma/2}u(t,\cdot)\right\|_{2}+\|u(t,\cdot)\|_{2}\right)^{p-1}+\left(\left\|\mathcal{L}^{\sigma/2}v(t,\cdot)\right\|_{2}+\|v(t,\cdot)\|_{2}\right)^{p-1}\right]$
$\times\left(\left\|\mathcal{L}^{\sigma/2}(u-v)(t,\cdot)\right\|_{2}+\|(u-v)(t,\cdot)\|_{2}\right).$
Now, using the above estimate for $v=0$ and the estimates from Theorem 2.5 for $u_{0}=0$ and $u_{1}=f(u)$, we get
$\|\partial_{t}^{i}\mathcal{L}^{j}u^{\operatorname{non}}(t)\|_{2}=\left\|\partial_{t}^{i}\mathcal{L}^{j}\int_{0}^{t}E_{1}(t-\tau)*_{\mathcal{L}}f(u)(t)\mathrm{d}\tau\right\|_{2}\leq\int_{0}^{t}e^{-\delta(t-\tau)}\left\|f(u)(\tau)\right\|_{2}\mathrm{d}\tau\\ \leq e^{-\delta t}\int_{0}^{t}e^{-\delta(p-1)\tau}\mathrm{d}\tau\|u\|^{p}_{X}\leq e^{-\delta t}\|u\|^{p}_{X}.$ (6.10)
@@
$\|\mathcal{N}u\|_{X}\leq A\|(u_{0},u_{1})\|_{\mathcal{A}^{\sigma,\theta}}+A\|u\|_{X}^{p},$ (6.11)
@@
Estimate for $\|\mathcal{N}u-\mathcal{N}v\|_{X}$: For $u,v\in X$, the Definition 6.6 yields
@@
Verifying that $\|u^{\mathrm{lin}}-v^{\mathrm{lin}}\|_{X}=0$ is straightforward. We must now estimate the following norms in order to calculate the norm $\left\|u^{\operatorname{non}}-v^{\operatorname{non}}\right\|_{X}$:
$\|\partial_{t}^{i}\mathcal{L}^{j}(u^{\operatorname{non}}-v^{\operatorname{non}})(t)\|_{2},\quad(i,j)\in\{(0,0),(0,\sigma/2),(1,0)\}.$
Utilizing the estimate analogous to the estimate $\left\|\partial_{t}^{i}\mathcal{L}^{j}u^{\operatorname{non}}(t,\cdot)\right\|_{2}$, we get
$\left\|\partial_{t}^{i}\mathcal{L}^{j}\left(u^{\operatorname{non}}-v^{\operatorname{non}}\right)(t)\right\|_{2}\lesssim\int_{0}^{t}e^{-\delta(t-\tau)}\|f(u)-f(v)\|_{2}d\tau\lesssim e^{-\delta t}\|u-v\|_{X}\left(\|u\|_{X}^{p-1}+\|v\|^{p-1}_{X}\right).$
@@
$\left\|\partial_{t}^{i}\mathcal{L}^{j}\left(u^{\operatorname{non}}-v^{\operatorname{non}}\right)(t)\right\|_{2}\lesssim\|u-v\|_{X}\left(\|u\|_{X}^{p-1}+\|v\|^{p-1}_{X}\right).$
Thus, combining the above estimate with $\|u^{\mathrm{lin}}-v^{\mathrm{lin}}\|_{X}=0$, we have
$\|\mathcal{N}u-\mathcal{N}v\|_{X}\leq B\|u-v\|_{X}\left(\|u\|_{X}^{p-1}+\|v\|^{p-1}_{X}\right),$ (6.12)
@@
Let us set $R:=rA(\|(u_{0},u_{1})\|_{\mathcal{A}^{\sigma,\theta}})$ for some $r>1$ with sufficiently small initial Cauchy data $\|(u_{0},u_{1})\|_{\mathcal{A}^{\sigma,\theta}}<\varepsilon$ such that
@@
$\|\mathcal{N}u\|_{X}\leq\frac{2R}{r}\quad\text{and }\quad\|\mathcal{N}u-\mathcal{N}v\|_{X}\leq\frac{1}{r}\|u-v\|_{X},$
respectively for all $u,v\in\mathscr{B}_{R}:=\{u\in X:\|u\|_{X}\leq R\}$. As a closed ball of Banach space $X$, $\mathscr{B}_{R}$ is a Banach space itself. This allows us to conclude the proof of the theorem using the Banach fixed theorem. ∎
@@ 7. Final remarks
    The presence of a discrete spectrum enables us to uniformly bound numerous terms, a feature that is generally not available in settings with a continuous spectrum, such as the classical wave equation or the wave equation on the Heisenberg group. This key difference facilitates the derivation of exponential decay estimates.
  2. (2)
@@
    The decay estimates in Theorem 2.5 remain valid for higher-order spatial and temporal derivatives, allowing the construction of global solutions with enhanced regularity compared to those in Theorem 6.3. In this higher-regularity framework, however, the admissible range of parameters becomes more constrained: in certain cases, both the spatial dimension (or, equivalently, the regularity index $\sigma$) and the nonlinearity exponent $p$ are restricted by a Gagliardo–Nirenberg inequality.
@@
    In Theorem 6.3, if we drop the assumption that $\mathcal{L}$ is strictly positive operator then we can only expect local well-posedness of solution due to lack of decay observed in Theorem 2.4. After local existence is established, the study naturally turns to long-term behavior, such as whether the solution may be extended worldwide or blows up finitely.